OpenStudy (anonymous):

how do you determine the range of a function? ex. a. f(x)=12x+3 b. g(x)=1/x^2 c. h(x)=square root of(5x-7) +4

OpenStudy (anonymous):

\[c. h(x)= \sqrt{5x-7}+4\]

OpenStudy (zzr0ck3r):

most of the time you can find the inverse of the function and look at its domain. y = 12x+3 (y-3)/12 = x domain is all R so range of 12x+3 is all real

OpenStudy (anonymous):

what do you mean by "domain is all R so range of 12x+3 is all real"? how did you get that

OpenStudy (zzr0ck3r):

the domain of (y-3)/12 is what?

OpenStudy (anonymous):

x?

OpenStudy (zzr0ck3r):

domain is numbers not variables. wht numbers can y be?

OpenStudy (zzr0ck3r):

well it can be a variable but dont think about that right now

OpenStudy (anonymous):

but as a number it can be anything

OpenStudy (zzr0ck3r):

yes, so we found the inverse function and its domain to be R so the originals functions range is all R

OpenStudy (zzr0ck3r):

I should say relation and not function

OpenStudy (anonymous):

so the range to any function with no fraction and no square root is all real numbers?

OpenStudy (zzr0ck3r):

no it could have a log most of the time the problem comes from 1/x sqrt(x) and ln(x) ... most of the time.

OpenStudy (anonymous):

ok anyway what about B how do you get the range when there's a fraction?

OpenStudy (zzr0ck3r):

I did that one, read up ^.

OpenStudy (zzr0ck3r):

y = 1/x^2 x^2 = 1/y x = +-sqrt (1/y) so 1/y must be >0 so y is all posotive numbers thus the range of 1/x^2 is all positive numbers